A patients temperature was 97.7 F at 6:10 A.M. and 98.9 at 1:10 P.M. If the temperature change with respect to elapsed time was linear throughout the day, what would the patient's temperature be at 7:10 P.M.?
Answer #1

Time passed between 6:10 am and 1:10 pm is, to = 7 hours

So, slope of the T-t curve, m = (98.9 - 97.7) / 7 = (6/35) F/hour

So, the equation for the T-t curve is,

T = 97.7 + (6/35)Δt

where Δt is the time elasped from 6:10 am.

So, temperature of the patient at 7:10 am is,

T = 97.7 + [(6/35) * 1] = 97.9 F